OpenStudy (18jonea):

Calculate the value of x in each of the following diagrams. Show all of your work

OpenStudy (18jonea):



OpenStudy (18jonea):

@jdoe0001

OpenStudy (jdoe0001):



OpenStudy (18jonea):

are you sure @jdoe0001

OpenStudy (18jonea):

so 88 +86 =174 / 2= 87